OpenStudy (anonymous):

What is the y intercept of x/7-4y=-2

OpenStudy (anonymous):

rearrange the equation into general form of straight line, y=mx+c which c is the y-intercept.

OpenStudy (anonymous):

How do you write this in standard form?

OpenStudy (anonymous):

x/7-4y=-2, -4y=-2- x/7, y=1/2 + x/28.

OpenStudy (anonymous):

I would put it in standard form by first multiplying both sides of the equation by 7 to get rid of the denominator in the first term: x/7 That would be \[(\frac{x}{7}-4y)*7 = -2*7\] then distributing on the left you get \[\frac{x}{7}*7-4y*7 = -2*7\] which is \[\frac{x}{7}*\frac{7}{1}-4y*7 = -2*7\] and the first term reduces to x since the 7's cancel so you get \[x-28y=-14\] Now isolate y on one side of the equation by putting all the terms without y on the other and dividing each term by -28

OpenStudy (anonymous):

\[x-28y=-14 \] -x -x \[\frac{-28y}{-28}=\frac{-x}{-28}-\frac{14}{-28}\] \[\frac{-28y}{-28}=\frac{-x}{-28}+\frac{-14}{-28}\] which simplifies to \[y=\frac{x}{28}+\frac{1}{2}\] c is \[\frac{1}{2}\] Does that help? :)